# Multi-Amplifier Sensing Charge-coupled Devices for Next Generation Spectroscopy

### v1

@@ 2 From Skipper to MAS CCDs
First developed as part of a U.S. Department of Energy Quantum Information Science initiative, the MAS CCD was designed to address the trade-off between reaching sub-electron levels of read noise and extended readout time that is a key drawback for single-amplifier Skipper CCDs (Holland, 2023). This is accomplished by implementing a chain of floating-gate output amplifiers, interspersed along an extended serial register. A schematic of this readout structure is illustrated in Fig. 1. Charge packets pass through and are sampled non-destructively by each output stage. Each charge packet is thus repeatedly measured to reduce noise in a similar approach to the Skipper CCD, but allows the charge to be continuously transferred to the next amplifier for the subsequent charge packet to be measured. This architecture permits sampling charge packets simultaneously for floating-gate amplifiers to effectively parallelize readout, thereby decreasing the readout time. Thus, the effective noise per pixel after averaging charge measurements for amplifiers is
@@ 2.1 Device properties
The test device was fabricated by Teledyne DALSA Semiconductor from high resistivity (10 k/cm) -type silicon substrate with buried -type channels and no backside processing or anti-reflective (AR) coating. There are 16 MAS readout channels along the 256-pixel extended serial register. On the opposite end of the serial register is a DESI-like output amplifier which, if connected, can enable the sensor to act as a conventional CCD when charge is clocked in the direction away from the MAS amplifiers. Standard three-phase clocking is employed to read out charge vertically and horizontally, and the serial clocks are similar to that of Skipper CCDs with the exception of one additional pixel separation (PS) gate. The PS gate acts as a potential barrier between floating-gate sense node and subsequent inter-amplifier pixels. This output structure is illustrated in the charge transfer diagram in Fig. 1. In the final output stage, a dump gate (DG) takes the place of the PS gate, and is structurally similar to the Skipper CCD output stage. Since the serial register is only one edge of this prototype detector, there is only one set of vertical clock lines, and the pads on the CCDs on both sides of the device that correspond to these vertical clocks are connected to the same set of clock drivers. The sets of serial clock voltages in the MAS register are also tied together, so all MAS output stages have the same set of clock voltages. The device has a format of 1024 x 512 pixels with 15 m pixels and since the sensor was not thinned, the thickness was 650 m and was illuminated on the frontside. These properties are summarized in Table 1.
@@ 2.2 Detector packaging
A “picture frame” format carrier was designed for the CCD with an opening slightly larger than the size of the sensor centered on a printed circuit board (PCB). The carrier board includes a 1 F and 4 M filter in parallel visible in Fig. 2, which helps stabilize the output voltage of the video transistor. A square silicon substrate was bonded to the back of the PCB, covering this opening. The backside of the CCD was then carefully bonded onto the substrate with epoxy through the opening on the front side of the PCB, then baked to cure. The CCD pads were then wirebonded with the PCB. A ribbon cable was used to connect this package to the dewar vacuum feedthrough. The wirebonded detector is pictured in Fig. 2.
@@ 3 Experimental apparatus and readout electronics
The dewar vacuum feedthrough links the ribbon cable from the CCD package to a dewar side board connected to the readout electronics. The CCD control is performed with an expanded version of the DESI front-end electronics (FEE) system to enable the simultaneous digital correlated double sampling (CDS) readout of the 16 output channels on the sensor. The new electronics developed, called the Hydra, runs as a set of four synchronized DESI FEE systems in a leader-follower configuration, consisting of a video processing board and power conversion board for each of the four FEE sub-units and a single clock driver board for the master FEE. The video board serves to process and digitize four channels of CCD video signal using a 100 MHz 16-bit analog-to-digital converter (ADC). The clock board provides the clock signals with a range 12 V for the CCD with a 40 channel 16-bit digital-to-analog converter (DAC). Each component of the FEEs within the Hydra sits on a motherboard with an FPGA which is the timing engine and reads ADC data output from the video board, and CDS data is sent to a MicroBlaze processor accessible by a TCP/IP communication interface. Data products are written out in standard FITS image format. Four main operating modes are defined for the CCD in our setup: exposing, readout, idle, and cleaning. Cleaning is performed prior to exposures and is intended to clear remnant charge from the active area by completing a readout sequence without reset or signal sampling. Idle mode continuously clocks the CCD between exposures and is stopped during a set exposure time and readout.
@@ 4 Noise performance
For a CCD image with zero exposure time, pixel values are the sum of the dark current accumulated during the readout time, the bias level DC offset, occasional cosmic ray signal, and amplifier read noise. As a functional definition, we refer to dark current in this work as the signal detected when the CCD is not exposed to light, and note that the thermal leakage current is only one component among other dominating components including stray light and thermal photons from the dewar walls; we do not attempt to separate these components in this study. Extended strips of virtual pixels along the image borders created by overclocking the CCD array comprise the overscan region, where the distribution of pixel values reflect the read noise and bias level. We note that these virtual pixels also contain other components such as the dark current and cosmic ray contributions accumulated during the readout of a rows and columns. Overscan subtraction for the entire image including the virtual pixels corrects for the bias shift of all pixels by taking the sigma-clipped mean of the overscan region and subtracting this from each pixel. During testing, we found this yields similar results to a line-by-line overscan correction where the sigma-clipped mean pixel value for each row of the serial overscan is subtracted from each physical and virtual pixel in the same row. We apply iterative sigma clipping which rejects outlier pixel charge contaminated by cosmic rays from the analysis. The standard deviation derived from fitting the resulting overscan pixel distribution after overscan subtraction and sigma clipping corresponds to the read noise.
@@
Because the MAS CCD produces the same number of images as the number of output amplifiers for a single sample per amplifier in an exposure, in our case 16, these image frames can be used to measure the noise of its corresponding amplifier. The 16 images are then averaged to measure the combined noise for the multiple charge measurements made per pixel. The effect of averaging on noise as a function of the number of MAS amplifiers is shown in Fig. 3 in comparison to the theoretical expectation from Eq. 1. However, before the signal is averaged, the frames must be regridded to account for the 16 inter-amplifier pixel shifts, where each image starting from the second amplifier in the MAS chain is shifted 16 pixels serially forwards compared to the image output from the preceding amplifier. This correction was implemented by shifting images from the second amplifier onwards to share the pixel grid of the image from the first amplifier. These steps of shifting, signal averaging, and overscan subtraction were applied to each image as part of our standard preprocessing.
@@ 4.1 Gain calibration
The detector gain is a conversion factor between counts or analog-digital units (ADU) from the ADC output and number of electrons the CCD recorded, which we measure with the commonly used technique of stimulating the CCD with 5.9 keV X-rays from the radioactive decay of a 55Fe source. Since the quantity of electron-hole pairs generated in a pixel per X-ray photon absorbed in the silicon bulk is proportional to the X-ray energy and known, the gain can be derived from the measured location of the 5.9 keV Mn K peak in the spectrum (Janesick, 2001). After obtaining five second exposures, the X-ray hits were extracted to build a histogram of the charge distribution. The Mn K X-ray peak was then fitted with a simple Gaussian to find the position of the peak in ADU. The conversion gain is obtained as the ratio of this peak position with 1620 e-, the average number of electrons produced by the K X-ray (Fraser et al., 1994). The gain for each amplifier is shown in Fig. 5 for our standard ADC gain setting and a reduced ADC gain used to extend the dynamic range to high signal levels ( electrons). Throughout the course of our tests in operating the CCD for one year, we have found the amplifier gains to be highly stable over time.
@@ 4.2 Voltages and noise optimization
The CCD was run with a substrate-bias voltage () of 40 V at which lateral charge diffusion is expected to be smaller than the pixel size (Karcher et al., 2003). Because we are interested in the performance at the fastest readout configuration where charges are sampled once for each amplification stage (i.e., no skipping), the output gate is biased to 0 V in our setup. The potential diagram with our voltages in the output structure is shown in Fig. 6, where each row of clock sequences represents a state during the readout. The output stage involves the summing gate (SG), output gate (OG), floating-gate (FG) sensing region, and pixel separation gate (PS) and are interleaved by three horizontal clock phases between each output stage. In the last stage, instead of a PS gate, a dump gate (DG) leads to , a bias voltage applied at -22 V to remove the charge from the channel. Because the sense node relies on a floating gate for non-destructive charge measurement, the exact potential here cannot be known and is therefore denoted as a dashed segment in Fig. 6. Several clock states are depicted in Fig. 6, where state (a) is where charge enters the sense node, state (b) is where the charge is being measured, and state (c) is where the charge exits the sense node and is shifted away to the next output stage in subsequent states.
@@
Since the prototype sensor is not divided into quadrants, we used a single set of three vertical clocks for the parallel register and three horizontal clocks for the serial register. Reducing the horizontal clock amplitudes was the most significant driver of noise reduction, allowing the noise floor to be lowered to near single-electron noise. Starting with a 15 V horizontal clock swing, the low voltages of the clocks were gradually increased as shown in Fig. 7. However, there is a limit to the minimum amplitude imposed by charge transfer efficiency between amplifier stages for the first horizontal clock phase, H1. This amplifier-to-amplifier charge transfer inefficiency is linked to the PS gate, where the low voltage of the PS gate must be at least 1 V greater than the H1 low state. When the H1 low state is within 1 V of the PS low state, the reset drain voltage () must be increased to allow potential for the PS low state to increase without causing inter-amplifier charge transfer streaking. Increasing the low voltage of the PS gate was crucial for raising the potential barrier between the sense node and the next stage of horizontal clock phases, when both the PS and H1 clocks are in low states, to prevent charge in one pixel from diffusing into a neighboring pixel. For instance, a PS gate voltage at -6.5 V provides a 5.5 V potential well for an H1 at -12 V whereas a PS gate voltage at -8 V only gives 4 V of potential barrier. In this latter case, the well depth formed by this voltage difference is smaller than the parallel clock amplitudes, which have a 5 V amplitude, and becomes a limiting factor for the effective well depth of the detector. At the minimum H1 amplitude with a high of 3 V and low of -8.5 V set globally for all horizontal clocks, the compound read noise reaches just under 1.1 e- rms/pixel with a single-amplifier read noise ranging from 3.68–4.51 e- rms/pixel depending on the amplifier. Although the returns diminish with this approach, it was possible to achieve 1.03 e- rms/pixel of compound noise without incurring charge transfer problems by tuning the second and third horizontal phases independently of the first phase. We interpret this noise to be an effective read noise because its tight correlation with serial clock amplitudes suggests that the noise includes not only intrinsic amplifier noise but also a serial clock-induced charge shot noise component. The serial clock-induced charge was found to a significant contributor to the measured noise at greater horizontal clock heights for a Skipper CCD with a 6000-pixel long serial register but reduces when these clock amplitudes are lowered (cf. Fig. 12 in Villalpando et al., 2024). Reducing the horizontal clock swings also reduces the full well capacity, which should be balanced against detector noise requirements, discussed in section 5.2.
@@
To estimate the cumulative effects of other noise components such as light leaks in the dewar and parallel register clock-induced charge, zero second exposure bias images were acquired to compare the noise in the overscan region and the active area. The virtual pixels of the overscan region are clocked only in the serial direction for each row. Consequently, differences in the pixel RMS between the two regions are expected to arise from the combination of light leakage and clock-induced charge. The bias images indicate a glow emanating from the top of the parallel register on the “left hand side” (on the prescan edge) of the device near a corner. The position of this glow directly opposite of where the serial register is located appears to rule out an amplifier origin. To prevent the signal from the glow from being transferred down the parallel register during idle mode, a stop idle-clear was implemented before readout. However, because the synchronization of the FEE subsystems requires the followers to be slightly ahead of the leader, there is a practical minimum time delay just under 0.5 s between when the clear is executed and the readout initiated. Thus, we restrict our noise calculations to the bottom 50 rows of the CCD near the serial register and only half the columns on the side towards the overscan where the glow signal has not yet been transferred in the 0.5 s of idle clocking time delay. We find that the difference in the standard deviation of the signal between the active area and overscan is about 0.34 e-. Since this noise quantity includes a combination of the glow, light leaks, and total parallel clock-induced charge, we interpret this difference in noise to be an upper limit on the clock-induced charge generated in the parallel register, demonstrating that this effect is not dominating our read noise measurements, making our noise calculation more robust.

### v2

@@ 2 From Skipper to MAS CCDs
First developed as part of a U.S. Department of Energy Quantum Information Science initiative, the MAS CCD was designed to address the trade-off between reaching sub-electron levels of read noise and extended readout time that is a key drawback for single-amplifier Skipper CCDs (Holland, 2023). This is accomplished by implementing a chain of floating-gate output amplifiers, interspersed along an extended serial register. A schematic of this readout structure is illustrated in Fig. 1. Charge packets pass through and are sampled non-destructively by each output stage. Each charge packet is thus repeatedly measured to reduce noise in a similar approach to the Skipper CCD, but allows the charge to be continuously transferred to the next amplifier for the subsequent charge packet to be measured. This architecture permits sampling charge packets simultaneously for floating-gate amplifiers to effectively parallelize readout, thereby decreasing the readout time. Thus, the effective noise per pixel after averaging uncorrelated charge measurements for amplifiers is
@@ 2.1 Device properties
The test device was fabricated by Teledyne DALSA Semiconductor from high resistivity (10 k-cm) -type silicon substrate with buried -type channels and no backside processing or anti-reflective (AR) coating. There are 16 MAS readout channels along the 256-pixel extended serial register. On the opposite end of the serial register is a DESI-like output amplifier which, if connected, can enable the sensor to act as a conventional CCD when charge is clocked in the direction away from the MAS amplifiers. Standard three-phase clocking is employed to read out charge vertically and horizontally, and the serial clocks are similar to that of Skipper CCDs with the exception of one additional pixel separation (PS) gate. The PS gate acts as a potential barrier between floating-gate sense node and subsequent inter-amplifier pixels. This output structure is illustrated in the charge transfer diagram in Fig. 1. In the final output stage, a dump gate (DG) takes the place of the PS gate, and is structurally similar to the Skipper CCD output stage. Since the serial register is only one edge of this prototype detector, there is only one set of vertical clock lines, and the pads on the CCDs on both sides of the device that correspond to these vertical clocks are connected to the same set of clock drivers. The sets of serial clock voltages in the MAS register are also tied together, so all MAS output stages have the same set of clock voltages. The device has a format of pixels with 15 m pixels and since the sensor was not thinned, the thickness was 650 m and was illuminated on the frontside. These properties are summarized in Table 1.
@@ 2.2 Detector packaging
A “picture frame” format carrier was designed for the CCD with an opening slightly larger than the size of the sensor centered on a printed circuit board (PCB). The carrier board includes a 1 F and 4 M filter in parallel visible in Fig. 2, which helps stabilize the output voltage of the video transistor by stabilizing the potential of the n+ guard structures (Holland et al., 2009). A square silicon substrate was bonded to the back of the PCB, covering this opening. The backside of the CCD was then carefully bonded onto the substrate with epoxy through the opening on the front side of the PCB, then baked to cure. The CCD pads were then wirebonded with the PCB. A ribbon cable was used to connect this package to the dewar vacuum feedthrough. The wirebonded detector is pictured in Fig. 2.
@@ 3 Experimental apparatus and readout electronics
The dewar vacuum feedthrough links the ribbon cable from the CCD package to a dewar side board connected to the readout electronics. The CCD control is performed with an expanded version of the DESI front-end electronics (FEE) system to enable the simultaneous digital correlated double sampling (CDS) readout of the 16 output channels on the sensor. The new electronics developed, called the Hydra, runs as a set of four synchronized DESI FEE systems in a leader-follower configuration, consisting of a video processing board and power conversion board for each of the four FEE sub-units and a single clock driver board for the master FEE. The video board serves to process and digitize four channels of CCD video signal using a 100 MHz 16-bit analog-to-digital converter (ADC). The clock board provides the clock signals with a range 12 V for the CCD with a 40 channel 16-bit digital-to-analog converter (DAC). Each component of the FEEs within the Hydra sits on a motherboard with an FPGA which is the timing engine and reads ADC data output from the video board, and CDS data is sent to a MicroBlaze processor accessible by a TCP/IP communication interface. Data products are written out in standard FITS image format. Four main operating modes are defined for the CCD in our setup: exposing, readout, idle, and cleaning. A clear is performed prior to exposures and is intended to clear remnant charge from the active area by completing a readout sequence without reset or signal sampling. Idle mode continuously clocks the CCD between exposures and is stopped during a set exposure time and readout.
@@ 4 Noise performance
For a CCD image with zero exposure time, pixel values are the sum of the dark current accumulated during the readout time, the bias level DC offset, occasional cosmic ray signal, and amplifier read noise. As a functional definition, we refer to dark current in this work as the signal detected when the CCD is not exposed to light, and note that the thermal leakage current is only one component among other dominating components including stray light and thermal photons from the dewar walls; we do not attempt to separate these components in this study. Extended strips of virtual pixels along the image borders created by overclocking the CCD array comprise the overscan region, where the distribution of pixel values reflect the read noise and bias level. A strip of 27 pixels called the prescan is also created along one edge of the CCD. These pixels are on the die with no associated column of active area pixels located between the first active column and the summing well of the first MAS output amplifier. We note that these pixels also contain other components such as the dark current and cosmic ray contributions accumulated during the readout of the rows and columns. Overscan subtraction for the entire image including the virtual pixels corrects for the bias shift of all pixels by taking the sigma-clipped mean111The sigma-clipped mean is an iterative process consisting in rejecting outliers (at standard deviation from the mean), and re-evaluating the mean and standard deviation of the remaining pixel values (e.g., astropy implementation). of the overscan region and subtracting this from each pixel. During testing, we found this yields similar results to a line-by-line overscan correction where the sigma-clipped mean pixel value for each row of the serial overscan is subtracted from each physical and virtual pixel in the same row. We apply iterative sigma clipping which rejects outlier pixel charge contaminated by cosmic rays from the analysis. The standard deviation derived from fitting the resulting overscan pixel distribution after overscan subtraction and sigma clipping corresponds to the read noise.
@@
Because the MAS CCD produces the same number of images as the number of output amplifiers for a single sample per amplifier in an exposure, in our case 16, these image frames can be used to measure the noise of its corresponding amplifier. The 16 images are then averaged to measure the combined noise for the multiple charge measurements made per pixel. The effect of averaging on noise as a function of the number of MAS amplifiers is shown in Fig. 3 in comparison to the theoretical expectation from Eq. 1. However, before the signal is averaged, the frames must be regridded to account for the 16 inter-amplifier pixel shifts, where each image starting from the second amplifier in the MAS chain is shifted 16 pixels serially forwards compared to the image output from the preceding amplifier. This correction was implemented by shifting images from the second amplifier onwards to share the pixel grid of the image from the first amplifier. These steps of shifting, signal averaging, and overscan subtraction were applied to each image as part of our standard preprocessing. An example of the 16-amplifier MAS CCD image output is shown in Fig. 4.
@@ 4.1 Gain calibration
The detector gain is a conversion factor between counts or analog-digital units (ADU) from the ADC output and number of electrons the CCD recorded, which we measure with the commonly used technique of stimulating the CCD with 5.9 keV X-rays from the radioactive decay of a 55Fe source. Since the quantity of electron-hole pairs generated in a pixel per X-ray photon absorbed in the silicon bulk is proportional to the X-ray energy and known, the gain can be derived from the measured location of the 5.9 keV Mn K peak in the spectrum (Janesick, 2001). After obtaining five second exposures where the front side of the CCD was illuminated by the 5.9 keV X-rays, the X-ray hits were extracted to build a histogram of the charge distribution. This detection of X-ray hits was performed with a standard source extraction algorithm which identifies sources with signal above the image background level (Barbary, 2016). This algorithm also performs conventional aperture photometry to extract the flux (e.g., Howell, 2006). Using a circular annulus with a minimum radius of two pixels and maximum outer radius of five pixels, the sum of the signal in the inner is subtracted by the background signal measured in the annular area between the inner and outer radii. With a histogram of the extracted fluxes, the Mn K X-ray peak was then fitted with a simple Gaussian to find the position of the peak in ADU. The conversion gain is obtained as the ratio of this peak position with 1620 e-, the average number of electrons produced by the K X-ray (Fraser et al., 1994). The gain for each amplifier is shown in Fig. 5 for our standard ADC gain setting and a reduced ADC gain used to extend the dynamic range to high signal levels ( electrons). Throughout the course of our tests in operating the CCD for one year, we have found the amplifier gains to be highly stable over time.
@@ 4.2 Voltages and noise optimization
The CCD was run with a substrate-bias voltage () of 40 V at which lateral charge diffusion is expected to be smaller than the pixel size. Because the CCD is front side illuminated, both photons from our light source and X-rays are converted to electron-hole pairs close to where the charge is stored and therefore diffusion is minimal (Karcher et al., 2003). Because we are interested in the performance at the fastest readout configuration where charges are sampled once for each amplification stage (i.e., no skipping), the output gate is biased to 0 V in our setup. The potential diagram with our voltages in the output structure is shown in Fig. 6, where each row of clock sequences represents a state during the readout. The output stage involves the summing gate (SG), output gate (OG), floating-gate (FG) sensing region, and pixel separation gate (PS) and are interleaved by three horizontal clock phases between each output stage. In the last stage, instead of a PS gate, a dump gate (DG) leads to the output drain with a voltage set to -22 V to remove the charge from the channel. Because the sense node relies on a floating gate for non-destructive charge measurement, the exact potential here cannot be known and is therefore denoted as a dashed segment in Fig. 6. Several clock states are depicted in Fig. 6, where state (a) is where charge enters the sense node, state (b) is where the charge is being measured, and state (c) is where the charge exits the sense node and is shifted away to the next output stage in subsequent states.
@@
Since the prototype sensor is not divided into quadrants, we used a single set of three vertical clocks for the parallel register and three horizontal clocks for the serial register. Reducing the horizontal clock amplitudes was the most significant driver of noise reduction, allowing the noise floor to be lowered to near single-electron noise. Starting with a 15 V horizontal clock swing, the low voltages of the clocks were gradually increased as shown in Fig. 7. However, there is a limit to the minimum amplitude imposed by charge transfer efficiency between amplifier stages for the first horizontal clock phase, H1. This amplifier-to-amplifier charge transfer inefficiency is linked to the PS gate, where the low voltage of the PS gate must be at least 1 V greater than the H1 low state. When the H1 low state is within 1 V of the PS low state, the reset drain voltage () must be increased to allow potential for the PS low state to increase without causing inter-amplifier charge transfer streaking. Increasing the low voltage of the PS gate was crucial for raising the potential barrier between the sense node and the next stage of horizontal clock phases, when both the PS and H1 clocks are in low states, to prevent charge in one pixel from diffusing back into the floating gate, an effect discussed in §5.1.2. For instance, a PS gate voltage at -6.5 V provides a 5.5 V potential well for an H1 at -12 V whereas a PS gate voltage at -8 V only gives 4 V of potential barrier. In this latter case, the well depth formed by this voltage difference is smaller than the parallel clock amplitudes, which have a 5 V amplitude, and becomes a limiting factor for the effective well depth of the detector. At the minimum H1 amplitude with a high of 3 V and low of -8.5 V set globally for all horizontal clocks, the compound read noise reaches just under 1.1 e- rms/pixel with a single-amplifier read noise ranging from 3.68–4.51 e- rms/pixel depending on the amplifier. Although the returns diminish with this approach, it was possible to achieve 1.03 e- rms/pixel of compound noise without incurring charge transfer problems by tuning the second and third horizontal phases independently of the first phase. We interpret this noise to be an effective read noise because its tight correlation with serial clock amplitudes suggests that the noise includes not only intrinsic amplifier noise but also a serial clock-induced charge shot noise component. The serial clock-induced charge was found to a significant contributor to the measured noise at greater horizontal clock heights for a Skipper CCD with a 6000-pixel long serial register but reduces when these clock amplitudes are lowered (cf. Fig. 12 in Villalpando et al., 2024). Reducing the horizontal clock swings also reduces the full well capacity, which should be balanced against detector noise requirements, discussed in section 5.2.
@@
To estimate the cumulative effects of other noise components such as light leaks in the dewar and parallel register clock-induced charge, zero second exposure bias images were acquired to compare the noise in the overscan region and the active area. The virtual pixels of the overscan region are clocked only in the serial direction for each row. Consequently, differences in the pixel RMS between the two regions are expected to arise from the combination of light leakage and clock-induced charge. The bias images indicate a glow emanating from the top of the parallel register on the “left hand side” of the device near a corner. The position of this glow directly opposite of where the serial register is located appears to rule out an amplifier origin. To prevent the signal from the glow from being transferred down the parallel register during idle mode where parallel clocks continuously shift charge down the pixel matrix when the CCD is not exposing or reading out, a stop idle to freeze clocking followed by a clear was implemented before starting the exposure. This clear action clears charge from the active area quickly by shifting out charge in the same clocking sequence as a readout but without reset or signal sampling. However, because the synchronization of the FEE subsystems requires the followers to be slightly ahead of the leader, there is a practical minimum time delay just under 0.5 s between when the clear is executed and the readout initiated. Thus, we restrict our noise calculations to the bottom 50 rows of the CCD near the serial register and only half the columns on the side towards the overscan where the glow signal has not yet been transferred in the 0.5 s of idle clocking time delay. We find that the difference in the standard deviation of the signal between the active area and overscan is about 0.34 e-. Since this noise quantity includes a combination of the light leaks and total parallel clock-induced charge, we interpret this difference in noise to be an upper limit on the clock-induced charge generated in the parallel register, demonstrating that this effect is not dominating our read noise measurements, making our noise calculation more robust.